OpenStudy (anonymous):

Can someone help me on this problem? Suppose part of the Diaz family goes in one balloon and other members go into another balloon. Because they fly at different altitudes and carry different weights, one balloon travels at 14 miles per hour and the other travels at 11 miles per hour. In how many hours will they be 12 miles apart?

OpenStudy (zzr0ck3r):

there is not way to answer this without more information

OpenStudy (zzr0ck3r):

if they take the exact same path, you can figure at one hour they will be 3 miles apart, at 2 they will be 6, and at 4 they will be ?

OpenStudy (anonymous):

well i did leave out some of the problem......"Each year, Alberquerque, New Mexico has a hot air balloon festival during which people can obtain rides in hot aIR BALLOONS.

OpenStudy (zzr0ck3r):

ok well they dont tell us if they are going the same direction. opposite direction.... but what I said in the last post still applies.

OpenStudy (anonymous):

AT 4 THEY WILL BE 12

OpenStudy (anonymous):

RIGHT?

OpenStudy (anonymous):

OH SORRY I LEFT OUT THAT THE OTHER BALLOON TRAVELS IN THE SAME DIRECTION AT 11 PER HOUR

OpenStudy (anonymous):

*MILES PER HOUR

OpenStudy (zzr0ck3r):

yes correct

OpenStudy (anonymous):

BUT COULD YOU SHOW ME HOW TO SET THE EQUATION UP PLEASE?

OpenStudy (zzr0ck3r):

um 14x -11x = 12

OpenStudy (zzr0ck3r):

d = r*t and so x is time for us and we have r*x-r*x = 12

OpenStudy (zzr0ck3r):

where 12 = d